OpenStudy (anonymous):

What is the fourth term in the expansion of (x + 2)6? 480x 168x4 160x3 120x3

OpenStudy (anonymous):

Fourth term will be: \[\huge T_4 = ^6C_3 \times (x)^3 \times (1)^3\] Can you find its value??

OpenStudy (anonymous):

whats t stand for and c

OpenStudy (anonymous):

T is for 4rth term and C stands for combination..

OpenStudy (anonymous):

the fourth term will be \(\dbinom{6}{3}x^3\times 2^3\) \[\dbinom{6}{6}=\frac{6\times 5\times 4}{3\times 2}=5\times 4=20\] \[2^3=8\] answer is \(8\times 20x^3=160x^3\)

OpenStudy (anonymous):

\[\huge ^nC_r = \frac{n!}{(n-r)! \times r!}\]